HUBBLE SPACE TELESCOPE DAILY REPORT #2969 PERIOD COVERED: 0000Z (UTC) 10/02/01 - 0000Z (UTC) 10/03/01 Daily Status Report as of 276/0000Z 1.0 OBSERVATIONS SCHEDULED AND ACCOMPLISHED: 1.1 Completed STIS/CCD 8903 (Bias Monitor - Part 1) The Space Telescope Imaging Spectrograph (CCD) was used to monitor the bias in the 1x1, 1x2, 2x1, and 2x2 bin settings at gain=1, and 1x1 at gain = 4 to build up high-S/N superbiases and track the evolution of hot columns. There were no problems. 1.2 Completed WF/PC-2 9042 (An Archive To Detect The Progenitors Of Massive, Core-Collapse Supernovae) The WF/PC-2 was used to search for supernovae which have massive star progenitors. The already extensive HST archive and high-resolution ground-based images of galaxies within ~20 Mpc enables us to resolve and quantify their individual bright stellar content. As massive, evolved stars are the most luminous single objects in a galaxy, the progenitors of core-collapse supernovae should be directly detectable on pre-explosion images. One Type II progenitor has been observed this year, and the investigators have proposed a short, companion WFPC2 proposal to confirm this candidate and identify a second. The observations completed nominally. 1.3 Completed WF/PC-2 8936 (Cycle 10 Supplemental Darks Pt1/3) The WF/PC-2 was used to perform a dark calibration program that obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The proposal completed with no reported problems. 1.4 Completed WF/PC-2 9057 (Host Galaxies of Obscured QSOs Identified by 2MASS) The WF/PC-2 was used to perform a snapshot survey of red QSOs discovered in The Two Micron All Sky Survey {2MASS} to investigate the detailed properties of their host galaxies. This large, possibly dominant, population of QSOs in the local universe has been previously overlooked because reddening by {intrinsic} obscuration along our line of sight causes their colors to be too red for identification by traditional "UV- excess" techniques. Their near-IR colors are similar to PG- type {UV-excess} QSOs, but it is far from certain whether they are indeed from the same parent population or represent a completely new class of QSO. There were no reported problems. 1.5 Completed Two Sets of STIS/CCD 9285 (POMS Test Proposal: STIS Non-scripted Parallel Proposal Continuation III) The Space Telescope Imaging Spectrograph (CCD) was used to make low galactic latitude, non-scripted parallel observations as part of a POMS test proposal. The observations were completed as planned, and no anomalies were reported. 1.6 Completed Two Sets of STIS/CCD 8905 (Read Noise Monitor) The Space Telescope Imaging Spectrograph (CCD) was used to measure the read noise of all the amplifiers {A, B, C, D} on the STIS CCD using pairs of bias frames. Full frame and binned observations are made in both Gain 1 and Gain 4, with binning factors of 1x1, 1x2, 2x1 and 2x2. All exposures are internals. The proposal completed with no reported anomalies. 1.7 Completed WF/PC-2/STIS/CCD 9267 (Supernova Search) The WF/PC-2 and the Space Telescope Imaging Spectrograph were used to search for high-redshift supernovae in GO-observed fields, taking advantage of good first-epoch observations and of the scheduling opportunities available because STIS is not currently observing. No anomalous activity was reported. 1.8 Completed Six Sets of WF/PC-2 8815 (Cycle 9 Earth Flats) The WF/PC-2 was used to monitor flatfield stability. This proposal obtains sequences of Earth streak flats to construct high quality flat fields for the WF/PC-2 filter set. These flat fields will allow mapping of the OTA illumination pattern and will be used in conjunction with previous internal and external flats to generate new pipeline superflats. The proposal completed nominally. 1.9 Completed STIS/CCD 8910 (Sparse Field CTE Test-Internal {Cycle 10}) The Space Telescope Imaging Spectrograph (CCD) was used to perform CTE measurements that are made using the "sparse field test", along both the serial and parallel axes. This program needs special commanding to provide {a} off-center MSM positionings of some slits, and {b} the ability to read out with any amplifier {A, B, C, or D}. No problems were encountered. 1.10 Completed Two Sets of WF/PC-2 8939 (Cycle 10 Internal Monitor) The WF/PC-2 was used to calibrate the internal monitor, to be run weekly to monitor the health of the cameras. As described in 2.1 and HSTAR 8360, the acquisition for the second iteration of this proposal defaulted to fine lock backup on one FGS only, possibly affecting the observations in this iteration. No other problems were encountered. 1.11 Completed Two Sets of STIS/CCD/MA1 9241 (The Disk and Wind of HD 104237) The Space Telescope Imaging Spectrograph (DDA and MA1) was used to perform GTO studies of intermediate-mass stars that have revealed circumstellar disks and associated nebulosities in 44% of our sample. The largest-scale nebulosity is seen in those systems with emission in the unidentified infrared bands, which have been interpreted as being associated with C-H stretch and bend modes in small organic grains. This hypothesis will be tested with coronagraphic observations of the nearby Herbig Ae star, HD 104237. As described in 2.1 and HSTAR 8360, the acquisition for the second iteration of this proposal defaulted to fine lock backup on one FGS only, possibly affecting the observations in this iteration. Otherwise, all other observations completed normally. 1.12 Completed Two Sets of WF/PC-2 9244 (POMS Test Proposal: WFII Parallel Archive Proposal Continuation) The WF/PC-2 was used to perform a generic target version of the Archival Pure Parallel program. The program will be used to take parallel images of random areas of the sky, following the recommendations of the Parallels Working Group. The observations completed with no anomalous activity. 1.13 Completed STIS/CCD 9051 (Identifying Damped Lyman-alpha Galaxies at z~1) The Space Telescope Imaging Spectrograph (CCD) was used to look for samped Lyman-alpha absorption systems that contain the bulk of the neutral gas in the Universe in the redshift range z = 0.5 - 5, yet the nature of the galaxies responsible for the absorption is not well understood. Only recently have observers found more than a handful of damped absorbers at redshifts z < 1.5. Using the FIRST Bright Quasar Survey {FBQS}, with over a 1000 quasars, the proposers have undertaken a survey to build a complete picture of he nature of the galaxies responsible for damped Lyman-alpha absorption systems at z~1 and to double the sample size at this redshift. No problems were encountered. 1.14 Completed STIS/CCD 9187 (A Lyman Limit Snapshot Survey: The Search for Candidate D/H Absorbers) The Space Telescope Imaging Spectrograph (CCD) was used to measure D/H in gas clouds seen against background QSOs. This provides a unique and fundamental cosmological probe of the baryonic density parameter. To date, most of the QSO absorption line D/H constraints have been derived at high redshift. In this proposal the proposers investigate a lower redshift range, because there are two potentially crucial advantages to be gained by carrying out D/H measurements at intermediate redshift. First, the background of HI forest lines is lower and contamination of the DI feature is significantly reduced. Second, a far larger sample of very bright QSOs exists. There were no reported problems. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions: 7 Successful: 7 Per HSTAR 8360, the acquisition at 275/165300Z defaulted to fine lock backup on FGS-3 only when the scan step limit was exceeded for FGS-1. The proposals detailed in 1.10 and 1.11 may have been affected. Scheduled Re-acquisitions: 7 Successful: 7 2.2 FHST Updates: Scheduled: 16 Successful: 15 The update at 276/002301Z failed; however, the following acquisition succeeded. 2.3 Operations Notes: The SSR EDAC error counter was cleared four times per ROP SR-1A. Using ROP DF-18A, the engineering status buffer limits were adjusted twice. The NSSC-1 status buffer was dumped at 276/0003Z as directed by ROP NS-03. A TTR was generated when there was a required re-transmit during a load uplink at 276/084308Z. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /CAW